# Thermodynamic and energetic constraints on out-of-equilibrium tunneling rates

### v1

@@ Abstract
We study bipartite quantum systems kept at different temperatures where a tunnel coupling between the two subsystems induces transitions. We find two independent constraints on the temperature-bias-dependent, out-of-equilibrium tunneling rates between the two subsystems, which both turn out to be particularly restrictive when the coupled quantum systems are small. These bounds take the form of a thermodynamic and of an energetic constraint, as they are associated with the dissipated heat and with the absorbed energy required to establish and deplete the temperature bias, respectively. The derived constraints apply to a large class of experimentally accessible quantum systems: except for the restriction to the tunneling regime, they hold for arbitrary subsystem Hamiltonians, including interactions or non-linear energy spectra. These results hold for a large class of experimentally relevant systems, ranging from molecular junctions to coupled cavities, and can be tested by, for instance, measuring the out-of-equilibrium tunneling current and its noise.
@@ I Introduction
Fluctuation theorems have been instrumental in studying the probability distribution of physical variables, such as thermodynamic work, in both classical and quantum stochastic thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In particular, detailed fluctuation theorems [7] constrain such probability distributions by relating the probability of a process to the probability of its time reverse. These relations provide a powerful framework to study fluctuations out of equilibrium, but they can also be used to, e.g., derive the equilibrium fluctuation-dissipation theorem (FDT) [17, 18, 19], which relates the fluctuations of observables to their dissipative responses. However, establishing a relation analogous to the FDT, linking generic correlations to response functions for systems out of equilibrium remains challenging [20, 21, 22]. For out-of-equilibrium correlated states, FDTs have been identified [23, 24, 25, 26, 27] for a generalized current operator, where its average and fluctuations are determined by two independent non-equilibrium transfer rates. Specifically, for a charge current induced by a voltage bias a FDT [28, 29] has been established far from equilibrium in the (weak) tunneling regime. In essence, this FDT extension relies on the detailed balance relation between these rates under the crucial assumption of a uniform temperature across the tunneling link. Consequently, the generalized FDT breaks down in the presence of a temperature bias.
@@
However, setups with a temperature bias are crucial in (quantum) thermodynamics, where they are used to fuel, e.g., heat engines [30, 31, 32]. Pivotal experiments have not only implemented nanoscale heat engines [33, 34, 35, 36] but also explored temperature biases for transport spectroscopy  [37, 38, 39, 40]. Importantly, systems exposed to large temperature biases also occur when one subsystem is cooled down with the help of coupling to another, possibly very different, subsystem [41, 42, 43]. Moreover, in nanoelectronics, systems with a temperature bias have attracted considerable attention in recent years in the context of so-called delta-$T$ noise [44]. Several experiments [45, 46, 47, 48] have detected this type of out-of-equilibrium charge noise generated by a temperature bias motivating further theoretical works [49, 50, 51, 52, 53, 54, 55, 56]. These diverse efforts call for a better understanding of how a temperature bias constrains the dynamics of small-scale devices.
@@
In particular, for nanoelectronics systems, fluctuation-dissipation bounds have recently been developed [57] for current fluctuations (or noise) in the presence of a temperature bias. While these bounds apply to conductors with generic transmission properties, their validity is limited to systems with weak electron-electron interactions. In the presence of possibly strong interactions, the perturbative approach developed in Refs. [24, 25, 26, 27] showed that noise is super-Poissonian in the tunneling regime, even in the presence of a temperature bias. However, while the approach is general and not restricted to electronic systems, this constraint is rather loose and the role of the temperature bias is not singled out. Hence, there remains the important question of whether fundamental bounds on the noise or, more generally, on the tunneling dynamics exist also for strongly interacting systems accounting for the impact of a temperature bias.
@@
In this work, we establish bounds on the nonequilibrium tunneling rates in the presence of a possibly large temperature bias accounting for the thermodynamic quantities required to generate such a temperature bias. To this end, we investigate the properties of the tunneling rates that characterize generic tunneling in a bipartite quantum system under arbitrary temperature biases. However, compared to earlier studies, we take a different approach in studying these out-of-equilibrium rates: Instead of comparing the rate of a process, say $\Gamma_{\rightarrow}$, with the rate of its reversed process, $\Gamma_{\leftarrow}$, we compare the full out-of-equilibrium rates for subsystems at different inverse temperatures, $\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})$, with the rates of the same process but in the absence of a temperature bias: $\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}})$. This approach allows us to prove two constraints on the tunneling rates which relate the cost of establishing a temperature bias to the response of the rates to a temperature change. These findings have direct implications on how the noise in systems subject to a temperature bias is constrained by the system dynamics. We thereby extend the scope of out-of-equilibrium noise at the intersection of quantum transport and quantum thermodynamics to systems with possibly strong interactions.
The main results of this paper are the thermodynamic constraint (Sec. III.1) and the energetic constraint (Sec. III.2) on the temperature-bias-dependent tunneling rates. These constraints connect the the out-of-equilibrium response of the tunneling rates to a temperature change to the thermodynamic and energetic costs of establishing a temperature bias, and take the form
$\mathcal{W}^{\text{(Thermo)}}\geq\mathcal{W}^{\text{(Resp)}},$ (1a)
$\mathcal{W}^{\text{(Energy)}}\geq\mathcal{W}^{\text{(Resp)}}.$ (1b)
Here,
$\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}\equiv\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})-\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),$ (2)
accounts for the response of $\Gamma_{\rightleftarrows}$ to a small change in the inverse temperature of the left (L) subsystem, both in the presence and in the absence of a temperature bias. Here, L is the subsystem that is taken at a different temperature when going from the configuration at a uniform temperature to that with a temperature bias. By contrast, $\mathcal{W}^{\text{(Thermo)}}$ and $\mathcal{W}^{\text{(Energy)}}$ are a thermodynamic and an energetic cost function, associated with the heat and with the energy required to change the temperature of subsystem L, respectively. More specifically, in the thermodynamic cost function
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}\equiv\widetilde{Q}^{\text{(c)}}\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}})+\widetilde{Q}^{\text{(h)}}\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),$ (3)
the tunneling rates with and without the temperature bias are weighted by the minimal amount of heat, $\widetilde{Q}^{\text{(h/c)}}$, dissipated to establish or to deplete the temperature bias of interest, $\beta_{\text{L}}-\beta_{\text{R}}$. In the energetic cost function,
$\mathcal{W}^{\text{(Energy)}}_{\rightleftarrows}\equiv\int_{U_{\text{L}}(\beta_{\text{R}})}^{U_{\text{L}}(\beta_{\text{L}})}\!\!\!\!\Gamma_{\rightleftarrows}(x,\beta_{\text{R}})d[U_{\text{L}}(x)]$ (4)
the tunneling rates are instead weighted by the amount of internal energy $dU_{\text{L}}(x)$ required for an infinitesimal increase of the temperature of subsystem L.
These general statements on the tunneling rates provide constraints also on the full out-of-equilibrium tunneling current noise, including the configuration with a temperature bias. This follows from a direct connection between the tunneling rates and both average and fluctuations of the tunneling current flowing between the two parts of the bipartite system. Importantly, our findings do not rely on any close-to-equilibrium fluctuation theorems and our results hold for generic bipartite quantum systems, where strong correlations within each subsystem might be present.
The remainder of this paper is organized as follows. After a brief description of our generic theoretical model in Sec. II, we present constraints on out-of-equilibrium rates in Sec. III. In Sec. IV, we showcase the constraints for two experimentally relevant example systems. Several Appendices provide details of our derivations of key equations.
II Theoretical model

### v2

# Thermodynamic and energetic constraints on transition probabilities of small-scale quantum systems

@@ Abstract
We study the transition probabilities of a two-point measurement on a quantum system, initially prepared in a thermal state. We find two independent constraints on the difference between transition probabilities when the system is prepared at different temperatures, which both turn out to be particularly restrictive when the measured quantum system is small. These bounds take the form of a thermodynamic and of an energetic constraint, as they are associated with the dissipated heat and with the absorbed energy required to increase or to reduce the temperature of the system. The derived constraints apply to arbitrary system Hamiltonians, including interactions or non-linear energy spectra. We show the relevance of these constraints for the special case where transitions are induced by energy or particle exchange in weakly coupled bipartite systems out of equilibrium. This example is of interest for a wide range of experimentally relevant systems, from molecular junctions to coupled cavities, and can be tested by, for instance, measuring the out-of-equilibrium tunneling current and its noise.
@@ I Introduction
Fluctuation theorems have been instrumental in studying the probability distribution of physical variables, such as thermodynamic work, in both classical and quantum stochastic thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In particular, detailed fluctuation theorems [7, 17] constrain such probability distributions by relating the probability of a process to the probability of its time reverse. These relations provide a powerful framework to study stochastic dynamics out of equilibrium, but they can also be used to, e.g., derive the equilibrium fluctuation-dissipation theorem (FDT) [18, 19, 20], which relates the fluctuations of observables to their dissipative responses. However, establishing a relation analogous to the FDT, linking generic correlations to response functions for systems out of equilibrium remains challenging [21, 22, 23]. For out-of-equilibrium correlated states, FDTs have been identified [24, 25, 26] for a generalized current operator, whose average and fluctuations are determined by two independent nonequilibrium transfer rates. Specifically, for a charge current induced by a voltage bias, a FDT [27, 28] has been established far from equilibrium in the (weak) tunneling regime. In essence, this FDT extension relies on the detailed balance relation between these rates under the crucial assumption of a uniform temperature across the tunneling link. Consequently, the generalized FDT breaks down in the presence of a more generic out-of-equilibrium situation, such as in the presence of a temperature bias.
@@
However, setups that can subsequently be in contact with environments at different temperatures, or even subject to a temperature bias are crucial in (quantum) thermodynamics, where they are used to fuel, e.g., heat engines [29, 30, 31]. Pivotal experiments have not only implemented nanoscale heat engines [32, 33, 34, 35] but also explored temperature biases for transport spectroscopy  [36, 37, 38, 39]. Importantly, systems exposed to large temperature biases also occur when one subsystem is cooled down with the help of a coupling to another, possibly very different, subsystem [40, 41, 42]. It is hence important to understand in which way coupling a system to different temperatures constrains its dynamics.
Figure 1: Two-point measurement schemes starting with an initially cold (a) or hot (b) state. The Gibbs probabilities (4) of observing the eigenstate of energy $\epsilon_{a}$ for a cold (blue) or hot (red) initial state (c) only cross at one energy, which we refer to as $\tilde{\epsilon}$.
In this work, we present general relations between transition probabilities in a two-point measurement scheme, comparing situations where the system is initially prepared in thermal states at different temperatures, see Fig. 1. This is a different approach than those in earlier studies on fluctuation relations [7, 15, 17], where transitions of a process are compared to transitions of the time-reversed process. Our approach allows us to establish bounds on the difference between transition probabilities: the main results of this paper are the thermodynamic constraint (Sec. II.2) and the energetic constraint (Sec. II.3) on the temperature-dependent transition probabilities. They relate the thermodynamic and energetic cost required to bring the system to these different temperatures with the response of the transition probabilities to a temperature variation. We do not make any assumptions on the Hamiltonian underlying the transition probabilities, which can hence involve strong interactions or nonlinear spectral properties.
One relevant application of our general findings are constraints on the dynamics of a bipartite system subject to a temperature bias. This could, e.g., be a small quantum system prepared at an initial temperature and coupled to an environment at a different temperature. As examples, we will choose bipartite systems with a weak tunnel coupling. The reason for this choice is the possibility to clearly identify the two different subsystem states and that the transition rates can be directly related to tunneling currents and the zero-frequency current noise [27, 28]. This allows us to exploit our findings to formulate constraints on noise in the spirit of FDTs but in the presence of a possibly large temperature bias, where standard FDTs are not applicable [18, 19, 20, 27, 28].
@@
Indeed, in particular for nanoelectronics systems, fluctuation-dissipation bounds have recently been developed [43] for current fluctuations (or noise) in the presence of a temperature bias. Although these bounds apply to conductors with generic transmission properties, their validity is limited to systems with weak electron-electron interactions. In the presence of possibly strong interactions, the perturbative approach developed in Refs. [25, 26] showed that noise is super-Poissonian in the tunneling regime, even in the presence of a temperature bias. However, the constraint is rather loose and the role of the temperature bias is not singled out. Hence, there remains the important question of whether fundamental bounds on the dynamics for strongly interacting systems exist accounting for the impact of a temperature bias.
@@
For such systems, using the derived general thermodynamic and energetic constraints, we establish bounds on the nonequilibrium tunneling rates in the presence of a possibly large temperature bias, accounting for the thermodynamic quantities required to generate such bias. These findings have direct implications on how the noise in temperature-biased systems is constrained by the system dynamics. We thereby extend the scope of out-of-equilibrium noise at the intersection of quantum transport and quantum thermodynamics to systems with possibly strong interactions. Importantly, our findings do not rely on any close-to-equilibrium fluctuation theorems.
The remainder of this paper is organized as follows. In Sec. II, we briefly introduce the concept of the two-point measurement with initial thermal states to then subsequently present a thermodynamic and an energetic constraints on the transition probabilities. In Sec. III, we apply our general results to experimentally relevant examples, which can be classified as weakly coupled bipartite systems in the presence of a large temperature bias. In Secs. III.3 and III.4, we showcase the constraints for an atom coupled to a cavity and two coupled fermionic tight-binding rings. Several Appendices provide details of our derivations of key equations.
II Constraints on dynamics with different initial states
Our goal is to compare the dynamics of two-point measurement schemes with different initial thermal states of the system and to constrain this difference by thermodynamic quantities. In this Section, we present general thermodynamic and energetic constraints.
II.1 Transition probabilities
We consider the general case in which a quantum system, initially prepared in the state described by the density matrix $\hat{\rho}$, undergoes a two-point projective measurement process. The first measurement is done on the basis {|$\{\ket{i}\}_{i}$, and has outcome $a$ with probability
$p_{a}=\braket{a}{\hat{\rho}}{a}.$ (1)
Then, the system undergoes an arbitrary unitary evolution $\hat{\mathcal{U}}(t,0)$ until time $t$, when a second measurement takes place. This last measurement is done on the basis {|$\{\ket{\psi_{i}}\}_{i}$, which may differ from the one of the initial measurement. The joint probability of measuring outcome $b$ in the second measurement, after the first measurement had outcome $a$, is given by
$p_{a\to b}=\left|\braket{\psi_{b}}{\hat{\mathcal{U}}(t,0)}{a}\right|^{2}p_{a},$ (2)
which is the probability of observing a transition $a\to b$ in the measurement outcomes.
These transition probabilities, compared with the transition probabilities in the time-reversed process are typically the starting point to develop fluctuation theorems [7, 15, 17]. Here, we are interested in finding out the impact of different temperatures on the dynamics of a system. We therefore start with the important statement that the temperatures only influence the initial state of the two-point measurement and not the conditional probabilities $|\braket{\psi_{b}}{\hat{\mathcal{U}}(t,0)}{a}|^{2}$. This means that we—instead of what is done in typical derivations of fluctuation theorems—need to compare the transition probabilities in the same process (i.e., induced by the same unitary evolution), but with different initial states. In particular, we consider initial states $\hat{\rho}$ being Gibbs states,
$\hat{\rho}=\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\,,$ (3)
where $\hat{H}_{0}$ is the system’s Hamiltonian at time $t=0$, $\beta\equiv T^{-1}$ is the inverse temperature of the system (note that temperatures have the units of energy, meaning that we set $k_{\text{B}}=1$), and $Z(\beta)\equiv\text{Tr}\left\{e^{-\beta\hat{H}_{0}}\right\}$ is the partition function. Furthermore, we take the first measurement to be in the energy eigenbasis $\hat{H}_{0}\ket{a}=\epsilon_{a}\ket{a}$, such that the probability of observing outcome $a$ in Eq. (1) reads
$p_{a}(\beta)=\frac{e^{-\beta\epsilon_{a}}}{Z(\beta)}$ (4)
where $\epsilon_{a}$ is the energy of |$\ket{a}$. We highlight the dependence on the inverse temperature $\beta$ by putting it as an argument of $p_{a}$ because, in the following, we compare initial states at different temperatures, as sketched in Fig. 1(a,b). Using properties of the Gibbs distribution, we establish constraints on the transition probabilities of Eq. (2) in the sections below that contain thermodynamic quantities, such as the internal energy $U$ and the entropy $S$. Those are given by
$U(\beta)$ $\equiv\text{Tr}\left\{\hat{H}_{0}\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\right\},$ (5a)
$S(\beta)$ $\equiv-\text{Tr}\left\{\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\log\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\right\}.$ (5b)
These quantities of the initial state are of interest for the constraints to be developed, since the temperature-dependence of the transition probabilities, Eq. (2), enters via the initial-state probability, Eq. (1), only.
II.2 Thermodynamic constraint
Figure 2: (a) An initially hot system is cooled down to $T_{\text{c}}$ using a cold bath. In this process, one can extract at most $-\Delta F^{\text{c}}$ [Eq. (9a)] as work. The now cold system undergoes the two-point measurement scheme. (b) An initially cold system is heated up to $T_{\text{h}}$ using a hot bath. In this process, one can extract at most $-\Delta F^{\text{h}}$ [Eq. (9b)] as work. The now hot system undergoes the two-point measurement scheme.
We start with an analysis of the initial state at different temperatures. When comparing a hot state at inverse temperature $\beta_{\text{h}}$ with a cold state at inverse temperature $\beta_{\text{c}}>\beta_{\text{h}}$, there is only one value of $\epsilon_{a}$ for which $p_{a}(\beta_{\text{h}})=p_{a}(\beta_{\text{c}})$, as shown in Fig. 1(c). This crossing energy is given by
$\tilde{\epsilon}=\frac{1}{\beta_{\text{c}}-\beta_{\text{h}}}\log\frac{Z(\beta_{\text{h}})}{Z(\beta_{\text{c}})}.$ (6)
Therefore the state-probabilities $p_{a}$ fulfill $p_{a}(\beta_{\text{h}})\gtrless p_{a}(\beta_{\text{c}})$ when $\epsilon_{a}\gtrless\tilde{\epsilon}$. It follows that
$[p_{a}(\beta_{\text{h}})-p_{a}(\beta_{\text{c}})][\epsilon_{a}-\tilde{\epsilon}]\geq 0$ (7)
for all $\epsilon_{a}$ and $\beta_{\text{c}}>\beta_{\text{h}}$. Summing (7) over all eigenstates $a$ leads to the statement that the internal energy is an increasing function of temperature, $U(\beta_{\text{h}})\geq U(\beta_{\text{c}})$. We furthermore note that the contributions of (7) can be rewritten in terms of a response of the probabilities to a temperature variation
$\partial_{\beta}p_{a}(\beta)=[U(\beta)-\epsilon_{a}]p_{a}(\beta).$ (8)
In contrast, the remaining term $[U(\beta_{\mathrm{h}})-\tilde{\epsilon}]p_{a}(\beta_{\text{h}})-[U(\beta_{\mathrm{c}})-\tilde{\epsilon}]p_{a}(\beta_{\text{c}})$ then includes thermodynamic quantities, in particular the nonequilibrium free energy differences, defined as
$\Delta F^{\text{c}}$ $\equiv[U(\beta_{\text{c}})-U(\beta_{\text{h}})]-T_{\text{c}}[S(\beta_{\text{c}})-S(\beta_{\text{h}})],$ (9a)
$\Delta F^{\text{h}}$ $\equiv[U(\beta_{\text{h}})-U(\beta_{\text{c}})]-T_{\text{h}}[S(\beta_{\text{h}})-S(\beta_{\text{c}})].$ (9b)
We find the relation
$[U(\beta_{\mathrm{h}})-\tilde{\epsilon}]p_{a}(\beta_{\text{h}})-[U(\beta_{\mathrm{c}})-\tilde{\epsilon}]p_{a}(\beta_{\text{c}})$ (10)
$=$ $-\frac{\beta_{\text{c}}\Delta F^{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}p_{a}(\beta_{\text{h}})-\frac{\beta_{\text{h}}\Delta F^{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}p_{a}(\beta_{\text{c}}).$
In this paper, we are interested in using the inequality (7) to formulate constraints for transition probabilities $p_{a\to b}$ in a two-point measurement, see Eq. (2). The above discussed insights can be readily transferred to transition probabilities, by simply multiplying (7), as well as (8) and (10), by the conditional probability $|\braket{\psi_{b}}{\mathcal{U}(t,0)}{a}|^{2}$, finding
$[p_{a\to b}(\beta_{\text{h}})-p_{a\to b}(\beta_{\text{c}})][\epsilon_{a}-\tilde{\epsilon}]\geq 0.$ (11)
Using the properties of (8) and (10) in (11), we establish the bound on transition probabilities
$\mathcal{W}_{a\to b}^{\text{Thermo}}\geq\mathcal{W}_{a\to b}^{\text{Resp}}\,,$ (12)
where we defined a temperature-response function
$\mathcal{W}_{a\to b}^{\text{Resp}}\equiv\partial_{\beta}p_{a\to b}(\beta_{\text{h}})-\partial_{\beta}p_{a\to b}(\beta_{\text{c}}),$ (13)
and a thermodynamic cost function
$\mathcal{W}_{a\to b}^{\text{Thermo}}\equiv-\frac{\beta_{\text{c}}\Delta F^{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}p_{a\to b}(\beta_{\text{h}})-\frac{\beta_{\text{h}}\Delta F^{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}p_{a\to b}(\beta_{\text{c}}).$ (14)
This thermodynamic constraint (12) on the transition probabilities at different temperatures is the first main result of this paper. It implies that the response of the transition probabilities to a change in temperature is limited by the transition probabilities themselves and by the thermodynamic cost of changing the system’s temperature. In order to intuitively understand how far it constrains the system dynamics at different initial temperatures, we consider the thermodynamics of cooling down a hot system in order to be able to subsequently perform the two-point measurement scheme starting from a cold state, as sketched in Fig. 2. Analogously, we consider the thermodynamics of heating up a cold system in order to be able to subsequently perform the two-point measurement scheme starting from a hot state, as sketched in Fig. 2.
  • (a)
    Cooling: The system initially at $T_{\text{h}}$ is cooled down by bringing it into contact with a bath at $T_{\text{c}}<T_{\text{h}}$ [Fig. 2(a)]. Heat flows out of the system until it reaches temperature $T_{\text{c}}$, thus inducing the change $[U(\beta_{\text{c}})-U(\beta_{\text{h}})]$ in internal energy and $[S(\beta_{\text{c}})-S(\beta_{\text{h}})]$ in entropy. While heat is flowing, it is possible to extract work, which is at most $-\Delta F^{\text{c}}$. After the system has been cooled down, the two-point measurement scheme is performed and leads to the transition probabilities $p_{a\to b}(\beta_{\text{c}})$.
  • (b)
    Heating: The system initially at $T_{\text{c}}$ is heated up by bringing it into contact with a bath at $T_{\text{h}}<T_{\text{c}}$ [Fig. 2(b)]. Heat flows into the system until it reaches temperature $T_{\text{h}}$, thus inducing the change $[U(\beta_{\text{h}})-U(\beta_{\text{c}})]$ in internal energy and $[S(\beta_{\text{h}})-S(\beta_{\text{c}})]$ in entropy. While heat is flowing, it is possible to extract work, which is at most $-\Delta F^{\text{h}}$. After the system has been heated up, the two-point measurement scheme is performed and leads to the transition probabilities $p_{a\to b}(\beta_{\text{h}})$.
Furthermore, the factors $\frac{\beta_{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{h}}}{T_{\text{h}}-T_{\text{c}}}$ and $\frac{\beta_{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{c}}}{T_{\text{h}}-T_{\text{c}}}$ in Eq. (14) correspond to the coefficient of performance of a heat pump and of a refrigerator, respectively. Consequently, the product $-\frac{\beta_{\text{c}}\Delta F^{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the heat absorbed by the cold bath during the cooling process when the extracted work is maximum, as detailed in Appendix A. Similarly, the product $-\frac{\beta_{\text{h}}\Delta F^{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the energy absorbed by the system during the heating process. Thus, in the thermodynamic cost function of Eq. (14) the hot and cold transition probabilities are weighted by the heat dissipated to cool down or heat up the system, respectively.
II.3 Energetic constraint
Figure 3: Heating stroke of the system. As the temperature increases from $T_{\mathrm{c}}$ to $T_{\mathrm{h}}$ (top to bottom panels), the energy absorbed, $dU$, and the transition probabilities, $p_{a\to b}$, are monitored at each intermediate temperature.
In order to set up the second constraint, instead of comparing the initial probability distribution, $p_{a}(\beta)$, at two different temperatures, we analyze how the distribution changes under an infinitesimal temperature variation. Differentiating Eq. (8), we have
$\partial_{\beta}^{2}p_{a}(\beta)=p_{a}(\beta)[U(\beta)-\epsilon_{a}]^{2}+p_{a}(\beta)\partial_{\beta}U(\beta).$ (15)
Since the first term on the right hand side is always positive, we can establish the inequality
$\int_{U(\beta_{\text{c}})}^{U(\beta_{\text{h}})}p_{a}(\beta)[dU(\beta)]\geq\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\ .$ (16)
Multiplying Eq. (16) by the conditional probability $|\braket{\psi_{b}}{\mathcal{U}(t,0)}{a}|^{2}$ reveals a constraint on the transition probability $p_{a\to b}$ of Eq. (2), given by
$\mathcal{W}_{a\to b}^{\text{Energy}}\geq\mathcal{W}_{a\to b}^{\text{Resp}}$ (17)
with the temperature-response function of Eq. (13) and the energetic cost function defined as
$\mathcal{W}_{a\to b}^{\text{Energy}}\equiv\int_{U(\beta_{\text{c}})}^{U(\beta_{\text{h}})}p_{a\to b}(\beta)[dU(\beta)].$ (18)
Equation (17) is the second central result of this paper. We refer to it as the energetic constraint since—unlike the thermodynamic constraint of Eq. (12)—it does not focus on the thermodynamic cost required to generate the temperature bias, but rather on the energetic cost. To understand the ingredients of this second constraint (17), one does not need to consider both heating and cooling processes depicted in Fig. 2, but only one continuous heating stroke as shown in Fig. 3. Starting from the system being at a cold temperature $T_{\mathrm{c}}=\beta_{\mathrm{c}}^{-1}$, we imagine slowly increasing its temperature to the hot temperature $T_{\mathrm{h}}=\beta_{\mathrm{h}}^{-1}$ and identifying the transition probabilities at each intermediate infinitesimal temperature change, as indicated by the interrupted yellow arrows. The energetic cost is found by weighting the transition probabilities with the corresponding variation of the internal energy needed to increase the temperature of the system.
II.4 Thermodynamic limit
Here, we comment on the relevance of the constraints (12) and (17) when the system approaches the thermodynamic limit. Notably, both the thermodynamic and the energetic constraints combine transition probabilities which are always smaller than one, with extensive properties of the system, namely with its internal energy and with its nonequilibrium free energy. However, the extensive quantities only appear on the left-hand side of Eqs. (12, 17), and not on the right-hand side. This feature implies that the left- and the right-hand sides behave very differently depending on the size of the system. More concretely, if the extensive quantities in the thermodynamic limit scale as
$U\to\lambda U,\qquad\Delta F^{\text{(c,h)}}\to\lambda\Delta F^{\text{(c,h)}}\ ,$ (19)
where $\lambda$ is the scale parameter, and the transition probabilities scale with an arbitrary scaling function $f(\lambda)$, i.e. $p_{a\to b}\to f(\lambda)p_{a\to b}$, both the thermodynamic constraint of Eq. (12) and the energetic constraint of Eq. (17) become trivial. Indeed, for $\lambda\to\infty$, the right-hand side containing the response $\mathcal{W}^{(\text{Resp})}_{a\to b}$ becomes negligible, and Eqs. (12, 17) then reduce to expressions that state the positivity of the thermodynamic and energetic costs, respectively, on the left-hand side.
However, the fact that the right hand sides of these equations can be neglected in the thermodynamic limit—independently on how it is taken—leads to a trivial statement. The thermodynamic and energetic costs are positive by construction: The transition probabilities are positive $p_{a\to b}\geq 0$, and so are the coefficients of performance $\eta^{\text{(c,h)}}\geq 0$, the nonequilibrium free energies $-\Delta F^{\text{(c,h)}}\geq 0$, as well as the energy variation $dU(x)$ in the integral in Eq. (17). Thus, both the thermodynamic and the energetic constraints pose constraints on small-scale quantum systems that do not satisfy the thermodynamic limit.
II.5 Saturating constraints and trivial constraints
The inequalities set up in Eq. (12) and Eq. (17) generically constrain the difference in transition probabilities in two-point measurement schemes for initial states at different temperatures. Here, we elaborate on the conditions under which these constraints are saturated and those under which they are trivial. By construction, both thermodynamic and energetic bounds become equalities at equal temperature, $\beta_{\mathrm{c}}=\beta_{\mathrm{h}}$. However, in this regime Eqs. (1217) become trivial since $\mathcal{W}^{\text{(Thermo)}}_{a\to b}=\mathcal{W}^{\text{(Energy)}}_{a\to b}=\mathcal{W}^{\text{(Resp)}}_{a\to b}=0$.
However, the constraints can be saturated non-trivially even for different temperatures when the energy of the initial state of the two-point measurement takes on specific values. The difference between Eq. (12) and Eq. (17) is also reflected in the conditions required to saturate the two constraints. For the thermodynamic constraint this energy corresponds to the energy at which the Gibbs distribution at different temperatures cross $p_{a}(\beta_{\mathrm{h}})=p_{a}(\beta_{\mathrm{c}})$. In this case, namely when $\epsilon_{a}=\tilde{\epsilon}$, the inequality (7) becomes an equality and hence (11) saturates. The energetic constraint can only be saturated if the internal energy of the system $U$ does not vary much in the considered interval, namely if $U(\beta_{\mathrm{c}})\approx U(\beta_{\mathrm{h}})$. If in addition, the energy of the initial state $\epsilon_{a}$ approximately equals this internal energy, the first term on the right hand side of Eq. (15) can be neglected and the energetic constraint (17) saturates.
In contrast, the constraints are always trivially fulfilled when the function characterizing the response to a temperature variation, $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ is negative. This is the case, when the energy of the initial state $\epsilon_{a}$ differs from the internal energy of the system by more than the internal energy’s standard deviation, see Appendix B. This means that the constraints are nontrivial only for those transition rates, where the initial state is “typical” for the system, namely with an energy close to the internal energy of the system.
III Particle and energy exchange in bipartite systems
Up to here, we have established and analyzed general constraints on transition probabilities in a two-point measurement scheme, without specifying how the transitions are induced. One topic of interest in which occurring transitions are detected is, e.g., in bipartite systems exchanging energy and particles. This could be a system coupled to an environment or more generally arbitrary coupled systems. In the following, we will apply the thermodynamic and energetic constraints developed in Sec. II on such a setting. For simplicity, we focus on a regime where the coupling between the two subsystems is weak. This has the advantage that a temperature bias between the two subsystems can be clearly defined. Furthermore, direct relations between transitions rates and fluctuating transport quantities can be established, thereby exploiting our results to pose constraints between currents and noise in the presence of a large temperature bias.
III.1 Weak tunnel coupling